# Breaking Performance Barriers: Massive-Scale Simulations of 2D Ising and Blume-Capel Models on Rack-Scale Multi-GPU Systems

### v1

@@ 1 Introduction
The two-dimensional (2D) Ising model has been simulated on various hardware platforms, including special-purpose machines with Field Programmable Gate Arrays (FPGAs) [1], Tensor Processing Units (TPUs) [2], the Cerebras Wafer-Scale Engine (WSE) [3], and Graphics Processing Units (GPUs) [4, 5, 6, 7, 8]. In this work, we update the highly optimized CUDA implementation of the 2D Ising model previously released in Ref. [4]. The two key innovations in this version are: $i)$ the ability to simulate much larger systems than in the previous version, and $ii)$ support for the Blume-Capel model in addition to the standard Ising model. Simulating a very large system, as opposed to performing multiple runs on smaller systems, offers two main advantages. First, it better matches the experimental scenario we aim to study: the growth of magnetic domains of size $\xi(t)$ in a system with linear size $L$ such that $L\gg\xi(t)$ (large enough that the experimental sample becomes effectively indistinguishable from an infinite system). Second, statistical errors scale with system size as $(\xi(t)/L)^{D/2}$, where $D$ is the spatial dimension ($D=2$ in our case). In other words, from a statistical perspective, a single simulation of a system with size $L^{\prime}=2L$ is equivalent to $2^{D}$ independent simulations of a system of size $L$.
@@ 3 Implementation
As the foundation for this work, we build upon the optimized GPU implementations presented in Ref. [4], which is a single-node/multi-GPU code that employs multi-spin coding (MSC). We also leverage CUDA Unified Memory to enable direct memory access between devices connected to the same machine via high-speed NVLink, eliminating the need for explicit data transfers.
@@ 3.1 Workload Distribution via Fabric Memory
As is common practice, the spin system is stored in memory using two separate buffers arranged in a checkerboard pattern. This partitioning ensures that all spins of one color can be updated in parallel. Since each spin’s flipping decision depends entirely on its neighbors in the opposite buffer, each simulation step consists of two update phases—one for each color. The workload is distributed across GPUs by partitioning each buffer into horizontal slabs, which are stored in device memory. This approach minimizes inter-GPU communication during spin updates, restricting it to just two rows: the last row of the slab on the preceding GPU and the first row of the slab on the next GPU. Notably, due to the periodic boundary conditions, the first and last GPU slabs are also connected. In our original code, this slab-based partitioning was easily implemented via Unified Memory. For each color, the host process simply called cudaMallocManaged once to request an allocation for the entire buffer. Then, cudaMemAdvise was called once per GPU to assign each slab (a specific address range) to its corresponding device’s global memory. This allows the GPUs to directly access any spin in the system by indexing the global buffers. The system automatically handles data transfers between devices by migrating memory pages as needed.
@@ 3.3 Correlation Protocol
$R=16,\quad r_{\rm c}(t)=\max\{256,\lfloor g(L)\sqrt{t}+0.5\rfloor\},\quad g(L)=6\sqrt{\frac{(\log(L/2^{16})}{3.3^{2}}+1.0}$
@@ 4 Results
Tables 1 and 2 present the single-GPU performance of the Ising and Blume-Capel codes on an NVIDIA GB200 GPU, varying the system’s linear size and covering total memory usage from $8$ MB to $128$ GB. For the Ising code, experiments were conducted with system sizes ranging from $L=2^{13}$ to $L=2^{20}$. Since each spin in the Blume-Capel code requires four bits instead of one, the runs for this case started at $L=2^{12}$ and went up to $L=2^{19}$. Peak performance for the Ising code is reached at $L=2^{16}$ ($1.0$ GB), exceeding $99\%$ of the absolute maximum performance of $1802$ updates/ns, which is achieved at $L=2^{20}$. For the Blume-Capel code, peak performance occurs at $L=2^{16}$, exceeding $99\%$ of the maximum performance of $948$ updates/ns, which is attained at $L=2^{19}$.
Table 3 summarizes the strong scaling measurements for both codes. For each model, we ran the same system using between $1$ and $64$ GPUs over $128$ time steps. The lattice sizes were selected to maximize memory usage on a single GPU, with $2^{38}$ spins ($L=2^{19}$) for the Blume-Capel model and $2^{40}$ spins ($L=2^{20}$) for the Ising model. Each experiment produced identical numerical results (within each model), regardless of the number of GPUs used. The figure above the table illustrates the corresponding speedup graph. The codes demonstrate near-perfect linear speedup, benefiting from the minimal inter-GPU communication required during the spin update step.
@@
Table 4 showcases the weak scaling measurements, where the system size per GPU is kept fixed while scaling from $1$ to $64$ GPUs. As in the strong scaling experiments, the lattice size per GPU was chosen to maximize memory usage. Specifically, for the Blume-Capel model, the system size per GPU was maintained at $2^{38}$ spins, while for the Ising model, it was set to $2^{40}$ spins. The figure above the table shows the corresponding efficiency plot. As in the strong scaling case, the flip update rate was measured over 128 simulation steps. Although increasing the number of GPUs results in larger system boundaries (two rows of $L$ spins) while keeping the number of spins per device constant, the relative ratio of boundary to total spins remains small. Consequently, access to remote GPU memory has no significant impact on performance.
@@
The highest spin update rates we measured with our codes are $114,729$ and $60,381$ updates/ns for the Ising and Blume-Capel implementations, respectively. To the best of our knowledge, the previous highest Ising flip rate was reported in Ref. [3], where the authors achieved $61,853$ updates/ns by running $754$ parallel simulations on a WSE, each with a size of $11586\times 16384$. Our results demonstrate that a GB200 NVL72 system—comparable to the WSE—delivers nearly twice the performance for Ising simulations (1.85$\sim$ 1.85$\times$ 1.85 ×). Additionally, the GB200 NVL72 enables the study of significantly larger systems (up to 370.000$\sim$ 370.000$\times$ 370.000 × larger), offering much higher statistical quality, since statistical errors scale inversely with system size in these simulations. For example, a single run on a system of size $L=2^{22}$ is free of finite-size artifacts, provided the coherence length remains below $\xi(t)\approx 2^{18}$. In Ref. [22], we simulated up to $\xi(t)\sim 2^{10}$, which would likely cause incipient finite-size artifacts in a system of linear dimensions $11586\times 16384$. Furthermore, the statistical errors for a single run of $L=2^{22}$ are roughly equivalent to those obtained from $2^{44}/(11586\times 16384)\approx 92675$ independent runs of a $11586\times 16384$ system.
@@ 5 Conclusions
In this study, we presented high-performance implementations of the two-dimensional Ising and Blume-Capel models, optimized for large-scale multi-GPU simulations. These implementations leverage multi-node NVLink systems, such as the NVIDIA GB200 NVL72. By employing advanced memory management techniques and optimizing Monte Carlo kernels, our implementations achieve unprecedented simulation scales, handling up to $2^{46}$ spins for the Ising model and $2^{44}$ spins for the Blume-Capel model. Our benchmarks demonstrate near-perfect strong and weak scaling up to 64 GPUs, highlighting the effectiveness of the NVLink interconnect in enabling memory sharing across multiple compute nodes. The Ising model achieves a peak update rate of $114,729$ updates/ns, significantly surpassing previous performance records. Meanwhile, the Blume-Capel model reaches $60,381$ updates/ns, marking the first high-performance implementation of this model to our knowledge. These results further demonstrate the versatility of GPUs as general-purpose accelerators, showing their ability to scale seamlessly to rack-level configurations. Advances in high-speed interconnects, unified memory architectures, and parallel programming frameworks now allow programmers to treat distributed GPUs as a unified compute resource.

### v2

# Massive-Scale Simulations of 2D Ising and Blume-Capel Models on Rack-Scale Multi-GPU Systems

@@ 1 Introduction
We present computer programs specifically developed to simulate the out-of-equilibrium dynamics of the two-dimensional (2D) Ising and Blume-Capel models—two paradigmatic systems that belong to the same equilibrium universality class [1, 2, 3]. Our chosen dynamics is the standard Metropolis spin-flip algorithm, see, e.g. [4, 5], representative of the so-called model-A dynamic universality class, which describes local dynamics without conservation laws [6]. Out-of-equilibrium dynamics is a central theme in Statistical Mechanics and plays a crucial role in the theoretical, numerical, and experimental study of disordered magnetic systems (see, e.g., [7, 8, 9, 10] for review). Importantly, it is also a vibrant area of research for non-disordered magnetic systems as well [11], with increasing interest in recent years (see, e.g., [12, 13, 14, 15, 16, 17]). It is important to emphasize that the physics studied here is specific to model-A dynamics. Consequently, simulation algorithms belonging to other dynamic universality classes—such as cluster methods [18, 19] or the Worm algorithm [20]—are not suitable for our purposes, despite their well-known efficiency in driving systems to thermal equilibrium.
In a typical experimental or computational setting, a system is quenched at the initial time $t=0$ from a high temperature down to a working temperature $T$, where it is then allowed to relax. The chosen temperature $T$ typically lies close to, or below, the critical temperature $T_{\mathrm{c}}$. The relaxation process involves the time-dependent growth of magnetic domains, characterized by a linear size $\xi(t)$. In a typical experiment (see, e.g., [10]), the domain size remains much smaller than the overall system size, i.e., $L\gg\xi(t)$.
To obtain results that are effectively indistinguishable from those in the large-$L$ limit, the system size should exceed the correlation length by at least a factor of (say) $50$. As a result, in numerical simulations, the system size effectively acts as a cutoff that limits the maximum correlation length $\xi(t)$ attainable during the simulation (because one needs to ensure that $L$ is at least as large as 50 $\xi(t)$ for all times considered in the simulation). In fact, simulating a single large system, rather than performing multiple runs on smaller systems, offers two significant advantages. First, it enables $\xi(t)$ to evolve over a broader range without being affected by finite-size effects. This is particularly important, as $\xi(t)$ serves as the fundamental scale variable in the scaling analyses central to the study of out-of-equilibrium dynamics (see, e.g., [17, 21, 22]). Second, statistical errors decrease with increasing system size, scaling as $[\xi(t)/L]^{D/2}$, where $D$ is the spatial dimensionality (with $D=2$ in our case). In practical terms, this means that a single simulation of a system with size $L^{\prime}=2L$ provides statistical precision equivalent to $2^{D}$ independent simulations of a system of size $L$.
The simplicity of spin models makes them particularly well-suited for implementation on dedicated hardware. Indeed, several computers have been purpose-built for the simulation of 2D and 3D spin systems. In parallel, highly specialized software has been developed for platforms such as Graphics Processing Units (GPUs). A portion of this work focuses on simulation algorithms that do not belong to the model-A dynamic universality class—for example, implementations on dedicated machines [23, 24, 25] and GPU-based approaches [26, 27, 28, 29, 30, 31]. However, considerable effort has also been invested in the implementation of model-A dynamics, both in hardware platforms [32, 33, 34, 35, 36, 37, 38, 39, 40] and in GPU-based software solutions [41].
Most directly related to our work are studies that apply model-A algorithms to simulations of the 2D ferromagnetic Ising model. This model has been simulated on various hardware platforms, including special-purpose machines with Field Programmable Gate Arrays (FPGAs) [42], Tensor Processing Units (TPUs) [43], the Cerebras Wafer-Scale Engine (WSE) [44], and Graphics Processing Units (GPUs) [45, 46, 47, 48, 49, 50]. In this work, we update the highly optimized CUDA implementation of the 2D Ising model previously released in Ref. [45]. The two key innovations in this version are: $i)$ the ability to simulate much larger systems than in the previous version (the importance of this point has already been emphasized above), and $ii)$ support for the Blume-Capel model in addition to the standard Ising model.
@@ 3 Implementation
As the foundation for this work, we build upon the optimized GPU implementations presented in Ref. [45], which is a single-node/multi-GPU code that employs multi-spin coding [64] (MSC). We also leverage CUDA Unified Memory to enable direct memory access between devices connected to the same machine via high-speed NVLink, eliminating the need for explicit data transfers.
@@ 3.1 Workload Distribution via Fabric Memory
As is common practice, the spin system is stored in memory using two separate buffers arranged in a checkerboard pattern.***The checkerboard lattice decomposition is crucial in numerical simulations and pure theoretical settings. For instance, the lattice decomposition in two sublattices played a significant role in [65]. This partitioning ensures that all spins of one color can be updated in parallel. Since each spin’s flipping decision depends entirely on its neighbors in the opposite buffer, each simulation step consists of two update phases—one for each color. The workload is distributed across GPUs according to a 1D decomposition, that is by partitioning each buffer into horizontal slabs, which are stored in device memory. This approach minimizes inter-GPU communication during spin updates, restricting it to just two rows: the last row of the slab on the preceding GPU and the first row of the slab on the next GPU. Notably, due to the periodic boundary conditions, the first and last GPU slabs are also connected. In our original code, this slab-based partitioning was easily implemented via Unified Memory. For each color, the host process simply called cudaMallocManaged once to request an allocation for the entire buffer. Then, cudaMemAdvise was called once per GPU to assign each slab (a specific address range) to its corresponding device’s global memory. This allows the GPUs to directly access any spin in the system by indexing the global buffers. The system automatically handles data transfers between devices by migrating memory pages as needed.
@@ 3.3 Correlation Protocol
$R=16,\quad r_{\rm c}(t)=\max\{256,\lfloor g(L)\sqrt{t}+0.5\rfloor\},\quad g(L)=6\sqrt{\frac{\log(L/2^{16})}{3.3^{2}}+1.0}$
@@ 4 Results
We assess the performance improvement of our current Monte Carlo implementation by comparing it to the previous version released in [45]. Both codes were executed on a single GB200 GPU, using the largest system size supported by the earlier implementation, $L=2^{19}$. That version stored $4$ bits per spin, requiring $128$ GB of memory for the full lattice. The previous code achieved a speed of $1070$ updates/ns, while the new implementation reaches 1800 updates/ns — an improvement of approximately 1.7×1.7$\times$1.7 ×.
Tables 1 and 2 present the single-GPU performance of the Ising and Blume-Capel codes on an NVIDIA GB200 GPU, varying the system’s linear size and covering total memory usage from $8$ MB to $128$ GB. The data from the tables are plotted in Figures 2 and 3. For the Ising code, experiments were conducted with system sizes ranging from $L=2^{13}$ to $L=2^{20}$. Since each spin in the Blume-Capel code requires four bits instead of one, the runs for this case started at $L=2^{12}$ and went up to $L=2^{19}$. Peak performance for the Ising code is reached at $L=2^{16}$ ($1.0$GB), exceeding $99\%$ of the absolute maximum performance of $1802$ updates/ns, which is achieved at $L=2^{20}$. For the Blume-Capel code, peak performance occurs at $L=2^{16}$, exceeding $99\%$ of the maximum performance of $948$ updates/ns, which is attained at $L=2^{19}$.
Table 3 summarizes the strong scaling measurements for both codes. For each model, we ran the same system using between $1$ and $64$ GPUs over $128$ time steps. The lattice sizes were selected to maximize memory usage on a single GPU, with $2^{38}$ spins ($L=2^{19}$) for the Blume-Capel model and $2^{40}$ spins ($L=2^{20}$) for the Ising model. Each experiment produced identical numerical results (within each model), regardless of the number of GPUs used. Figure 4 illustrates the corresponding speedup graph. The codes demonstrate near-perfect linear speedup, benefiting from the minimal inter-GPU communication required during the spin update step (each GPU exchanges $0.25$ MB and $0.5$ MB of data with its two neighbors for the Ising and Blume-Capel models, respectively).
@@
Figure 2: Plot showing the spin update throughput data from Table 1.
@@
Figure 3: Plot showing the spin update throughput data from Table 2.
Figure 4: Plot showing the strong scaling performance of the Ising and Blume-Capel codes, based on the data from Table 3.
@@
Table 4 showcases the weak scaling measurements, where the system size per GPU is kept fixed while scaling from $1$ to $64$ GPUs. As in the strong scaling experiments, the lattice size per GPU was chosen to maximize memory usage. Specifically, for the Blume-Capel model, the system size per GPU was maintained at $2^{38}$ spins, while for the Ising model, it was set to $2^{40}$ spins. Figure 5 shows the corresponding efficiency plot. As in the strong scaling case, the flip update rate was measured over 128 simulation steps. Although increasing the number of GPUs results in larger system boundaries (two rows of $L$ spins) while keeping the number of spins per device constant, the relative ratio of boundary to total spins remains small (with 64 GPUs, each device exchanges $2$ MB and $4$ MB of data with its two neighbors for the Ising and Blume-Capel models, respectively). Consequently, access to remote GPU memory has no significant impact on performance.
Figure 5: Plot showing the weak scaling performance of the Ising and Blume-Capel codes, based on the data from Table 4.
@@
The highest spin update rates we measured with our codes are $114729$ and $60381$ updates/ns for the Ising and Blume-Capel implementations, respectively. To the best of our knowledge, the previous highest Ising flip rate was reported in Ref. [44], where the authors achieved $61853$ updates/ns by running $754$ parallel simulations on a WSE, each with a size of $11586\times 16384$. Our results demonstrate that a GB200 NVL72 system—comparable to the WSE—delivers nearly twice the performance for Ising simulations (1.85$\sim$ 1.85$\times$ 1.85 ×). Additionally, the GB200 NVL72 enables the study of significantly larger systems (up to 370000$\sim$ 370000$\times$ 370000 × larger), offering much higher statistical quality, since statistical errors scale inversely with system size in these simulations. For example, a single run on a system of size $L=2^{22}$ is free of finite-size artifacts, provided the coherence length remains below $\xi(t)\approx 2^{18}$. In Ref. [17], we simulated up to $\xi(t)\sim 2^{10}$, which would likely cause incipient finite-size artifacts in a system of linear dimensions $11586\times 16384$. Furthermore, the statistical errors for a single run of $L=2^{22}$ are roughly equivalent to those obtained from $2^{44}/(11586\times 16384)\approx 92675$ independent runs of a $11586\times 16384$ system.
@@ 5 Conclusions
In this study, we presented high-performance implementations of the two-dimensional Ising and Blume-Capel models, optimized for large-scale multi-GPU simulations. These implementations leverage multi-node NVLink systems, such as the NVIDIA GB200 NVL72. By employing advanced memory management techniques and optimizing Monte Carlo kernels, our implementations achieve unprecedented simulation scales, handling up to $2^{46}$ spins for the Ising model and $2^{44}$ spins for the Blume-Capel model. Our benchmarks demonstrate near-perfect strong and weak scaling up to 64 GPUs, highlighting the effectiveness of the NVLink interconnect in enabling memory sharing across multiple compute nodes. The Ising model achieves a peak update rate of $114729$ updates/ns, significantly surpassing previous performance records. Meanwhile, the Blume-Capel model reaches $60381$ updates/ns, marking the first high-performance implementation of this model to our knowledge. These results further demonstrate the versatility of GPUs as general-purpose accelerators, showing their ability to scale seamlessly to rack-level configurations. Advances in high-speed interconnects, unified memory architectures, and parallel programming frameworks now allow programmers to treat distributed GPUs as a unified compute resource.
A natural direction for future developments would be to extend the codes to support 3D systems. Although the shared buffer based on fabric memory and the optimized random bit generation could be reused with minimal modifications, the spin processing kernels would likely require substantial changes. A 3D lattice may benefit from a different spin layout in memory, which would necessitate a new memory access pattern for reading neighboring spins. Furthermore, careful consideration should be given to the domain decomposition strategy to minimize intra-GPU communications when computing correlations.